Fuller Company has identified 4 different activity pools for its overhead costs, allocated costs to each pool and estimated activity levels as summarized in the following tables: Estimated activity levels for the year Pool 1 Number of setups 600 Pool 2 Inspection hours 2000 Pool 3 Machine hours 28000 Pool 4 Number of parts 1800 Direct labor hours 30800 Direct materials cost 144000 Pool 1 Pool 2 Pool 3 Pool 4 Setups Quality control Fabricating Ordering Total estimated costs in each pool $68400 $70000 $336000 $45000 1. What is the activity rate for pool 1? 2. What is the activity rate for pool 3? 3. Assume Fuller used a single plant-wide rate based on direct labor hours. What would that rate be? (Round the rate to the nearest dollar.)

Alphabet Company has identified 4 different activity pools for its overhead costs. They have completed an allocation of all costs to the pools other than supervisory costs, which are estimated at $68000 for the year. The controller has now determined that the supervisory salaries are allocable to the 4 pools in the following manner:

Setups Inspections Grinding Receiving Total
Supervisory Salaries 22% 5% 36% 37% 100%



The controller has also estimated the following activity levels for each pool:

Estimated activity levels for the year
Pool 1 Number of setups 1500
Pool 2 Inspection hours 4500
Pool 3 Machine hours 38000
Pool 4 Number of parts 31500


In the table below, complete the calculation of total costs allocable to each pool (stage 1 allocation). Then compute the activity rate for each pool. Round any answer to the nearest dollar.

Pool 1 Pool 2 Pool 3 Pool 4
Setups Inspections Grinding Receiving
Supervisory Salaries $ $ $ $
All other costs $79540 $131600 $203520 $100840
Total costs $ $ $ $
Activity rate: $ $ $ $
Answer #1

1.What is the activity rate for pool 1 = Overhead Cost/Activity Level

= 68,400/600

= $114 per Set up

2.What is the activity rate for pool 3

= 336,000/28,000

= $12 per Machine Hour

3.Single Rate = Total Overheads/Total Direct labor Hours

= 519,400/30,800

= $16.86 per hour i.e. $17 per labor hour

Pool 1

Pool 2

Pool 3

Pool 4

Setups

Inspections

Grinding

Receiving

Supervisory Salaries

$14,960

$3,400

$24,480

$25,160

All other costs

$79540

$131600

$203520

$100840

Total costs

$94,500

$135,000

$228,000

$126,000

Activity rate:

$63

$30

$6

$4

Note: Supervisory Salaries have been allocated in the ratio given.

Activity rate has been calculated by dividing total cost by the total quantity of driver